OpenStudy (anonymous):

Which function below has the largest slope? f(x) x f(x) -3 -10 -1 -7 1 -4 g(x) g(x) equals one-half x plus 3 h(x) Lance gains 1 level each time he plays his favorite video game. He started at level 10. f(x) g(x) h(x) The functions all have the same slope

OpenStudy (anonymous):

@saigealexander can you tell me what the slope of g(x) is?

OpenStudy (anonymous):

I have no clue @billj5

OpenStudy (anonymous):

the slope of f(x) = (y2-y1)/(x2-x1)=( -7 -(-10))/(-1 -(-3))=3/2 'g(x) equals one-half x plus 3 h(x)' there is any info about h(x). so i am assuming it miss type so it would be come like this g(x) equals one-half plus 3 f(x) this it could be solve g(x)=x/2+3 f(x) y-yo=m(x-xo) y-(-10)=3/2(x-(-3)) y+10=3/2x+9/2 y=3/2x+9/2-10=3/2x-11/2 f(x)=y=3/2x-11/2 g(x)=x/2+3(3/2x-11/2) g(x)=x/2+9/2x-33/2 g(x)=10/2 x -33/2 g(x)=5x -33/2....... the slope of g(x)=5